LBTO UT Apr 19: INAF Service and SO Training

(All times are MST)

Night crew:

Mountain: David G-H

Tucson: Michelle, Barry, Jenny, Shane, Andrew

Summary

LBTO Service Observing for INAF. Started the night with high priority Guaita in good seeing (0.8"-0.9") which very quickly degraded into very poor seeing 2.0-3.0" and recovered to variable 1.5-2.0" seeing.

Attempted Set 1 of low priority Smart program but ran into many problems with poor scripting and preparation. Worked through this got set and comparison stars complete.

Moved on to De Barros MOS program in better (0.9") seeing. Successfully completed 3 sets of observations and Tellurics

Finished with half of H2 imaging data for Target 2 of Beltran and accompanying telluric.

Sunset Calibrations

18:25 David is opening the enclosure.

18:30 Configuring for LUCI2 NB NIR skyflats. Telescope fixed at zenith.

18:35 Guider is on axis.

18:40 Counts in Br_gam are still a little high, ~30k.

18:48 Waiting for levels to drop to a reasonable level. Still too high.

18:54 Almost there, looks like we\x92ll have the first good one at sunset.

18:57 Sunset. First good (~10k) Br-gamma flats.

Br_gam NB LUCI2 Flats

NDIT=1

20160419.0015 10k

20160419.0016 10k

20160419.0017 10k

20160419.0018 10k

20160419.0019 9k

20160419.0020 8k

20160419.0021 8k

20160419.0022 8k

20160419.0023 7k

20160419.0024 6.5k



H2 NB LUCI2 Flats

NDIT=2

20160419.0026 5.8k

NDIT=5

20160419.0027 5k

NDIT=10

20160419.0028

We were in normal rather than integrated mode! H2 file numbers thus far are not correct.

19:06 DIT=3.0, NDIT=5, integrated, 20160419.0064 11k

20160419.0065 10.7k

20160419.0066 9.5k

20160419.0067 8.6k

20160419.0068 7.8k

20160419.0069 7.1k

20160419.0070 6.8k

20160419.0071 6.3k

20160419.0072 5.8k

20160419.0073 5k

19:10 Switching to Ks, putting telescope in track.

Ks broadband Flats

NDIT=10, DIT=3.0, integrated

20160419.0074 20k

NDIT=5, DIT=3.0

20160419.0075 8k

NDIT=10, DIT=3.0

20160419.0076 17k

20160419.0077 17k

20160419.0078 17k

20160419.0079 17k

20160419.0080 16.8k

20160419.0081 16.8k

20160419.0082 16.8k

20160419.0083 16.8k

20160419.0084 16.8k

20160419.0085 16.8k

Science Log

19:30 Slewing to standard star for Gauita Standard P550-C

19:34 Shell RIP

19:46 Pointing issues. Not on target yet.

19:47 12 deg twilight.

19:53 Initial DIMM seeing estimate: 0.8\x94. Collimating now.

19:56 0.7\x94 on guider. Slewing to Guaita Standard P550-C

19:58 Collimating. 0.8\x94 on guider, AM 1.18.

20:00 ~ 0.8\x94 on Ks image. Massive seeing bubble. Guider now 3\x94, 2.8\x94 on DIMM. Seeing bubble\x85 1.6\x94 on science image.

20:05 Waiting out the seeing bubble.



20:09 Seeing is still grim. We are considering our options but the only possibility is the backup sent to us late this morning (Smart). Searching for appropriate target

20:15 DIMM 2.4\x94, guider 1.8\x94

20:25 Backup program has significant issues with the scripts.

20:35 LUCI MOS issue. Called DThompson; grabber went beyond the rotation tolerance. Only needed an init, but none of us had seen this error before.

21:00 LUCI issue resolved, however we now have another pointing problem.

21:05 Acquiring Set 1 Telluric 3 for Smart; Seeing remains at >1.5. Skies are very clear.

21:16 Acquisition Set 1 Telluric 3 for Smart looks good; starting spectra

21:18 Another scripting issue detected. Script says Hi-res grating and z+J spec filter, checking with Felice

21:33 Script observing Set 1 Telluric 3

21:38 Slewing to Set 1 Science Target for Smart

21:45 Acq images 129-131. Acq image scripts use z+J; fixing.

21:54 Image 132; Acquisition confirmed

21:55 Set 1 Science Target science started

22:02 Slewing to Smart Set 1 Control 1

22:12 Guide star is too faint and seeing too poor, skipping to Smart Set 1 Control 3. Pointing problem; putting mask into FPU while problem is resolved

22:28 Instrument state in scheduler didn\x92t match instrument because of mask in FPU. Scheduler balked; restarted. IT 5814

22:31 Back in business. Slewing to Smart Set 1 Control 3

22:32 Spoke too soon, scheduler weirdness is transpiring (still IT 5814)

22:38 Ok, now we\x92re back in business

22:40 Script issues

22:54 Fixed script, reslewing; Observing Smart Set 1 Telluric 3

23:22 Slewing to next target, Smart Set 1 Telluric 3

23:50 Scheduler problems again.

23:55 Slewing back to previous control target with updated guide star: Smart Set 1 Control 1

00:20 Slewing to next target, Smart Set 9 Telluric 3

Seeing has been improving considerably over the last 30 minutes. We are switching to higher priority programs after this is complete. In total, we completed 1 program target from Set 1 and its telluric and 2 required control stars and their tellurics for Smart.



00:33 Slewing to next target, De Barros Telluric. The seeing has been stable under 1.3 for the last ~20 mins and continues to improve.

01:00 Slewing and acquiring MOS DeBarros Science

01:27 Acquisition done; starting science. First Sequence De Barros Science

02:04 Second Sequence De Barrow Science. Seeing is down to 0\x94.65 on the guider

02:34 Starting the Third Sequence De Barrow Science. Seeing is 0\x94.65.

In good seeing and excellent transparency, we finished 3 Sequences of De Barros high priority MOS mask and its accompanying telluric.

03:08 Switching to imaging Beltran Target 2, H2 imaging. Seeing is 0.8\x94

03:12 Pointing check

03:23 Images are elongated. At some dithers (larger off-source offsets) the guider is vignetted and may be contributing. Seeing is still good (0.8\x94)

03:52 Imaging Beltran Target 2 H2 Imaging continuing. Seeing = 0.7\x94 Transparency = clear

04:10 Seeing has worsened, over 1.3\x94 These last few images 270; 269 may be suspect.

04:15 Seeing definitely worse. Slewing to Standard 1 for Beltran H2 Imaging

04:18 Pointing check again. Resending Standard 1 for Beltran H2 Imaging

04:25 Guide star is too faint, switching to alternative standard target Standard 2 for Beltran H2 Imaging

04:40 Guide star resolved in first standard. Finishing with Standard 1 for Beltran H2 Imaging

Complete Beltran Target 2 H2 Filters in clear skies and good to variable seeing. Seeing started out quite good (0.7-0.8) but softened over course of last 15-20 minutes of science target, so PIs may want to check images.

04:50 12 deg twilight.

Sunrise Calibrations

05:02 Ks Band Flats

1 X (10 X 3) Nexpo X (NDIT X DIT)

tracking at zenith with random small dithers

0282 (23k) ; 0283 (23k); 0284 (23k); 0285 (23k); 0286 (23k); 0287 (26k); 0288 (25k); 0289 (28k); 0290 (30k); 0291 (45K); 0292 (60k); 0293 (80k)

05:30 Close

05:41 Sunrise

-- %USERSIG{MichelleEdwards - 2016-04-19}%

Comments

 
Topic revision: r2 - 21 Apr 2016, MichelleEdwards
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback